OpenStudy (anonymous):

-6.8-(2/5+3/10). help please

OpenStudy (anonymous):

-7.5

OpenStudy (anonymous):

first solve the bracket (2/5+3/10)=7/10 -6.8-(7/10)

OpenStudy (anonymous):

7/10=0.7 -6.8-0.7=-7.5

OpenStudy (anonymous):

no problem